OpenStudy (anonymous):

|(z)-(z/3)|=1 find the range to which |z| belongs

OpenStudy (anonymous):

z is complex? If so z=x+iy \[|x+iy-\frac{x+iy}{3}|=1\]

OpenStudy (anonymous):

\[|\frac{2x+2iy}{3}|=1\] \[\frac{2x+2iy}{3}\] Is just some other complex number with real part\[\frac{2}{3}\] and imaginary part also\[\frac{2}{3}\]

OpenStudy (anonymous):

 Where w is that complex number. Find the modulus by pythagoras, and you have an equation with a, b and constants.

OpenStudy (anonymous):

\[|w|=\sqrt{(\frac{2x}{3})^2+(\frac{2y}{3})^2}=1\]